# MedicoSAM: Towards foundation models for medical image segmentation

### v1

@@ Abstract
Medical image segmentation is an important analysis task in clinical practice and research. Deep learning has massively advanced the field, but current approaches are mostly based on models trained for a specific task. Training such models or adapting them to a new condition is costly due to the need for (manually) labeled data. The emergence of vision foundation models, especially Segment Anything, offers a path to universal segmentation for medical images, overcoming these issues. Here, we study how to improve Segment Anything for medical images by comparing different finetuning strategies on a large and diverse dataset. We evaluate the finetuned models on a wide range of interactive and (automatic) semantic segmentation tasks. We find that the performance can be clearly improved for interactive segmentation. However, semantic segmentation does not benefit from pretraining on medical images. Our best model, MedicoSAM, is publicly available. We show that it is compatible with existing tools for data annotation and believe that it will be of great practical value.
@@ 1 Introduction
Foundation models are large deep neural networks, often based on the transformer architecture [vaswani2017attention], which are trained on diverse datasets, either using self-supervised or supervised training objectives. They learn powerful representations that enable different downstream tasks either through in-context learning or finetuning. They underlie recent advances in language processing [brown2020language] and are also gaining importance in computer vision, thanks to the vision transformer (ViT) [dosovitskiy2020image]. The first foundation model that has gained wide-spread adoption for image segmentation is the Segment Anything Model (SAM) [kirillov2023segment]. It was trained on a large dataset of natural images with object annotations, using a supervised training objective that mimics interactive annotation. The model supports interactive and automatic segmentation tasks and generalizes to many different imaging modalities. More recently, SAM2 [ravi2024sam] has extended SAM to video data through architectural changes and a large video dataset with objects tracked over time. It supports interactive video segmentation in different modalities.
@@
SAM has been widely studied by the medical imaging community. Initial work has evaluated it for medical segmentation tasks (e.g. [mazurowski2023segment, ji2024segment, he2023computer]). The model showed impressive performance given that it was predominantly trained on natural images, but could not yet compete with domain specific models, especially for difficult tasks such as spine segmentation in MRI [mazurowski2023segment], segmentation of small organs in CT [huang2024segment], and other examples [cheng2023sam, he2023computer]. Consequently, follow-up work has improved SAM for medical images, either by finetuning it for interactive segmentation [ma2024segment, cheng2023sam, wang2023yanzhou, li2024polyp] or by using it as a pretrained encoder for semantic segmentation [zhang2023customized, wu2023medical, chen2023ma, gu2024build, yue2024surgicalsam, bui2023sam3d]. The model has also been adapted in related domains, for example to improve segmentation in microscopy [archit2023segment] and histopathology [horst2024cellvit]. Some work [ma2024segment, cheng2023sam, gu2024build] also tried to build a better foundation model for medical images, by finetuning SAM on a large medical dataset and publishing the updated model weights. Preliminary studies have also investigated SAM2 for medical images [dong2024segment, zhu2024medical, ma2024segment2]. They found mixed results for 2D segmentation, improving over SAM for some modalities but with worse performance for others, and promising results for video segmentation [liu2024surgical].
@@
However, a comprehensive study that compares different approaches for improving SAM as a foundation model for medical images is so far missing. In particular, prior work has not yet considered different criteria such a model should fulfill together:
@@
    It should improve interactive segmentation. The imaging modalities and segmentation tasks in medicine are very diverse. Hence, a single model that can solve any medical segmentation task fully automatically is currently not feasible [^1] There exist some efforts to establish such models for specific modalities, most notably TotalSegmentator [wasserthal2023totalsegmentator, d2024totalsegmentator] for CT and MRI images.. Better interactive segmentation will enable semi-automatic data annotation, leading to faster annotation times, either for data analysis or model training.
@@
    It should improve the performance for downstream tasks, in particular as a pretrained encoder for semantic segmentation. This would enable fully automating segmentation tasks by supervised finetuning; potentially using annotations generated interactively via (1).
@@
    It should be compatible with the original SAM library and graphical data annotation tools that use it (e.g. [liu2023samm, archit2023segment]), enabling users to benefit from improved interactive segmentation (1).
Previous work has only addressed one of these aspects at a time: MedSAM [ma2024segment] and SAMed2D [cheng2023sam] only evaluate interactive segmentation (1), Gu et al. [gu2024build] only study semantic segmentation (2). They do not explicitly study use of their models in user-friendly tools (3), and we find that only MedSAM is compatible with them.
Figure 1: a) Contribution overview: We finetune SAM on a large medical dataset to build MedicoSAM. We evaluate it for interactive segmentation and semantic segmentation, which requires training on additional annotated data (that could be generated via interactive segmentation), for 2D and 3D data. b) Results for interactive 2D segmentation, comparing MedicoSAM and other models derived from SAM. We report the average over 16 datasets for initial segmentation with a point (green) or box (yellow) and segmentation after iterative correction starting from a point (dark green) or box (dark purple). c) Results for interactive 3D segmentation. We report the average over 6 different datasets for segmentation based on a single point or box. d, e). Results for semantic 2D and 3D segmentation. We report the average over 6 datasets in both cases. The three best methods are highlighted, with MedicoSAM indicated by a dashed outline if it is not among the best models.
Our work closes this gap by comparing different training approaches on the dataset published by SAMed2d [cheng2023sam] and evaluating their efficacy for (1-3). Specifically, we compare different interactive-style finetuning approaches using our own implementation as well as extant SAM-based models. We compare these models for challenging medical segmentation tasks from four different categories: (i) interactive 2D segmentation, (ii) interactive 3D segmentation, (iii) semantic 2D segmentation, and (iv) semantic 3D segmentation. Where applicable, we also compare to SAM2, though we do not finetune it for medical data due to the recent release of this model and the lack of support in user-friendly tools. An overview of our approach is shown in Fig. 1 a) and the summary of our evaluation results is shown in Fig. 1 b-e). Overall, we find that domain specific finetuning can clearly improve interactive segmentation in 2D and 3D, given the right training objective. For semantic segmentation, finetuning does not provide a clear advantage as models pretrained on medical images do not consistently improve over the original SAM model. We also find that SAM-based architectures for semantic segmentation are not consistently better than nnU-Net [isensee2021nnu]. Our code and our best model, which we call MedicoSAM, are available at https://github.com/computational-cell-analytics/medico-sam.
@@ 2 Methods
We provide a summary of the contributions made by SAM, focusing on its training objective (Sec. 2.1). We then describe the finetuning approaches used in our study (Sec. 2.2), the extension of SAM to interactive 3D segmentation (Sec. 2.3), the methods for 2D and 3D semantic segmentation (Sec. 2.4), and our evaluation methodology (Sec. 2.5). We use these methods to compare different strategies to improve SAM for medical images and to train and publish MedicoSAM.
2.1 Segment Anything
@@
SAM [kirillov2023segment] is a vision foundation model for segmentation tasks. It consists of the image encoder, a ViT [dosovitskiy2020image], the prompt encoder and the mask decoder. This architecture enables the model to solve interactive segmentation tasks based on user input, so called prompts. The image encoder processes the image and outputs an image representation. It contains the majority of parameters. SAM provides three different versions with different encoder sizes, ViT-Huge (ViT-h), ViT-Large (ViT-l) and ViT-Base (ViT-b). The prompt encoder processes the prompts, which can be point coordinates, either a positive point prompt (within the object of interest) or a negative point prompt (outside of the object of interest), a box coordinate, or a low-resolution mask. It outputs a representation of the prompts; point, box and mask prompts can be combined. The mask decoder processes the outputs of image encoder and prompt encoder to predict a mask of the object of interest and a score that estimates the prediction quality. It has two heads. One predicts a single mask and score, the other predicts three masks and scores. The second head is for the case of a single point prompt, which can result in ambiguities for part-object segmentation. Fig. 2 a) shows an overview of SAM’s architecture.
@@
SAM was evaluated for a wide range of segmentation tasks in diverse image modalities, where it showed remarkable generalization. These capabilities are mainly due to two factors: the large and diverse training set and the sophisticated training objective. The training dataset, which is called SA-1B, consists of eleven million images with one billion annotated objects. It was generated by human annotators using SAM for semi-automatic annotation, followed by retraining and further annotation with the updated model. This procedure was repeated multiple times. The model was trained on this dataset using a supervised training objective that mimics interactive object annotation and correction: For a given ground-truth mask, the objective first samples either a point or box prompt and then corrects the model predictions with point prompts in multiple steps. In each step it computes $L_{mask}$, the loss between true and predicted mask as well as $L_{iou}$, the loss between the intersection over union (IOU) of true and predicted mask and the predicted score. These losses are accumulated and averaged at the end of a training iteration. See Alg. 1 for the pseudo-code of a training iteration. Besides this objective, the training proceeds as usual for deep neural networks by updating model weights with a version of stochastic gradient descent over multiple epochs.
Input: Images and ground-truth annotations, hyperparameters $n_{obj}$, $n_{steps}$, $p_{box}$, $p_{mask}$
@@
1 Sample minibatch of images and ground-truth
@@ 9             Compute bounding box of $m$, add as prompt to $p$
12             Sample random point from $m$, add as as positive point prompt to $p$
@@ 23       Compute IOU $i$ between $\hat{m}$ and $m$
24       Compute regression loss $L_{iou}(\hat{i}$, $i$)
@@ 27             Sample positive point from $m\,\&\,!\hat{m}$, add to $p$
28             Sample negative point from $!m\,\&\,\hat{m}$, add to $p$
@@ 34            Run lines 12-21 with current $p$
37Average losses in $L$, perform backprop
@@ 38 Update model parameters via optimizer
Algorithm 1 The training objective of SAM. This pseudo-code describes the open-source implementation from [archit2023segment], which slightly differs from [kirillov2023segment].
Recently, SAM was extended to videos by SAM2 [ravi2024sam]. This works extends the architecture with a memory bank to store prompts and mask predictions from previous frames, enabling interactive video segmentation. SAM2 was trained on a large annotated video dataset. While SAM2 is promising in the medical domain to analyze videos or volumetric data, it has so far received fewer attention. Initial studies [ma2024segment2, dong2024segment, zhu2024medical] have evaluated it for this domain and also performed preliminary transfer learning experiments, but so far have not provided an improved SAM2 model for medical data. Here, we include SAM2 in the evaluation of interactive segmentation, but do not finetune the model on medical data due to its recency and missing support in tools for medical data annotation.
@@ 2.2 Finetuning Segment Anything
We implement and compare different approaches for finetuning SAM to improve it for medical images. These approaches start by assembling a large labeled dataset of medical images and finetuning SAM on it using a variation of Alg. 1. The training algorithm has not been published by [kirillov2023segment], so each publication has used a custom implementation. The finetuning strategies also differ in which model parts they update, and whether an adapter-based strategy like LoRA [hu2021lora] is used.
@@
The authors of MedSAM [ma2024segment] assemble a dataset of ca. 1.5 million masks from CT, Endoscopy, MRI, X-Ray and other modalities, based on previously published data. They introduce a simple training objective that uses only box prompts, which are derived during training from the mask annotations. This objective corresponds to setting $n_{steps}=0,\,p_{box}=1,\,p_{mask}=0$ in Alg. 1. They use SAM with ViT-b encoder and update all parameters of the image encoder and mask decoder, freezing parameters of the prompt encoder. The finetuned model is evaluated for interactive segmentation based on box prompts. The model is publicly available.
@@
The authors of SAM-Med2D [cheng2023sam] build the SA-Med2D-20M dataset [ye2023sa], which consists of ca. 20 million masks in ca. 5 million images. The data covers ten different modalities (CT, Endoscopy, MRI, ultrasound, X-Ray and others) and is collected from published data. They finetune using the full training objective with $n_{steps}=8,\,p_{box}=0.5,\,p_{mask}=1$. However, they deviate from Alg. 1 in two ways: they sample either 1, 3, 5 or 9 points per step instead of a single positive and negative one (cf. lines 23-24) and they do not compute gradients for the prompt encoder in the steps corresponding to lines 22-29. They use ViT-b and update the parameters of all model parts. They use a custom adapter for the image encoder, similar to [wu2023medical], where a low rank projection layer is inserted between attention and feed-forward layer of each transformer block. Only the parameters of this layer are updated during training, the others are frozen. In addition, they train the model with a smaller input image size of 256 x 256 pixels instead of 1024 x 1024 pixels used by SAM. They also study a simpler finetuning strategy called FT-SAM where only the mask decoder is updated during training, image and prompt encoder are frozen. The finetuned models are evaluated for interactive segmentation and are publicly available.
@@
The authors of [gu2024build] assemble a dataset for finetuning comprising ca. 100,000 masks in ca. 300,000 images from CT, MRI, X-Ray and ultrasound. This dataset also contains unlabeled images that are used for self-supervised training. They compare two different finetuning objectives: updating the image encoder in a self-supervised manner using MAE [he2021mae] and using a simple supervised strategy with a box or a single point prompt, corresponding to $n_{steps}=0,\,p_{box}=0.5,\,p_{mask}=0$ in Alg. 1. They further compare finetuning the different model sizes with and without the use of LoRA [hu2021lora]. They evaluate these models for semantic segmentation downstream tasks. None of the models are publicly available.
In summary, prior work has studied different finetuning objectives and different model update strategies. To study the influence of the objectives we build on the flexible implementation of Alg. 1 provided by $\mu$SAM [archit2023segment], which was developed for microscopy data. We extend this implementation to also support simpler schemes (single box prompt / single box and point prompt) within the same framework. We do not study self-supervised training and finetune all model parts without the use of adapter layers. The first choice is due to the fact that self-supervised training would very likely lead to a loss of interactive segmentation performance. The second choice because we want to provide models that are compatible with the SAM library and tools using it. Introducing adapter layers would make the model incompatible and thus not practically useful, see also Sec. 3.3. Consequently, we study three different finetuning strategies:
  • MedSAM* (adapted from [ma2024segment]) uses only a box prompt, corresponding to $n_{steps}=0,\,p_{box}=1,\,p_{mask}=0$ in Alg. 1.
  • SimpleFT* (adapted from [gu2024build]) uses a single box or a single point prompt, corresponding to $n_{steps}=0,\,p_{box}=0.5,\,p_{mask}=0$ in Alg. 1.
@@
The parameter $n_{obj}$ is set to five for all three cases and we use the Dice loss for $L_{mask}$ and the L2 loss for $L_{iou}$. We finetune these three models on SA-Med2D-20M [ye2023sa]. We also benchmark the published models MedSAM [ma2024segment], SAM-Med2D [cheng2023sam], FT-SAM [cheng2023sam], SAM [kirillov2023segment], and where applicable SAM2 [ravi2024sam]. We use ViT-b for all models.
Figure 2: a) The SAM architecture for interactive segmentation consists of image encoder, prompt encoder (split into a part for mask prompts and a part for point/box prompts) and mask decoder. b) Adaptation for semantic segmentation following [zhang2023customized] (2D) and [chen2023ma] (3D). User prompts are replaced with a learnable prompt. The mask decoder is repurposed to predict the semantic segmentation. For 3D segmentation patch tokens across the depth axis are processed by additional 3D convolutions inserted in the image encoder.

### v2

# MedicoSAM: Robust Improvement of SAM for Medical Imaging

@@ Abstract
Medical image segmentation is an important analysis task in clinical practice and research. Deep learning has massively advanced the field, but current approaches are mostly based on models trained for a specific task. Training such models or adapting them to a new condition is costly due to the need for labeled data. The emergence of vision foundation models, especially Segment Anything Model (SAM), offers a path to universal segmentation for medical images, overcoming these issues. Here, we study how to improve SAM for medical images by comparing different finetuning strategies on a large and diverse dataset. We evaluate the finetuned models on a wide range of interactive and automatic semantic segmentation tasks. We find that performance clearly improves given the correct choice of finetuning strategies. This improvement is especially pronounced for interactive segmentation. Semantic segmentation also benefits, but the advantage over traditional segmentation approaches is inconsistent. Our best model, MedicoSAM, is publicly available. We show that it is compatible with existing tools for data annotation and believe that it will be of great practical value.
@@ 1 Introduction
Foundation models are large deep neural networks, often based on the transformer architecture [67], trained on diverse datasets, either with a self-supervised or supervised objective. They learn powerful representations that enable different downstream tasks either through in-context learning or finetuning. They underlay recent advances in language processing [4] and are also gaining importance in computer vision, thanks to the vision transformer (ViT) [13]. The first foundation model that has gained wide-spread adoption for image segmentation is the Segment Anything Model (SAM) [33]. It was trained on a large dataset of natural images with object annotations, using a supervised training objective that mimics interactive annotation. The model supports interactive and automatic segmentation tasks and generalizes to many different imaging modalities. More recently, SAM2 [61] has extended SAM to video data through architectural changes and a large video dataset with objects tracked over time. It supports interactive video segmentation in different modalities.
@@
SAM has been widely studied by the medical imaging community. Initial work has evaluated it for medical segmentation tasks (e.g. [55, 30, 22, 80, 39]). The model showed impressive performance given that it was predominantly trained on natural images, but could not yet compete with domain specific models, especially for difficult tasks such as spine segmentation in MRI [55], segmentation of small organs in CT [26], and other examples [9, 22]. Consequently, follow-up work has improved SAM for medical images, either by finetuning it for interactive segmentation [49, 9, 68, 43] or by using it as a pretrained encoder for semantic segmentation [79, 70, 8, 18, 77, 5]. The model has also been adapted in related domains, for example to improve segmentation in microscopy [1] and histopathology [24, 17]. Some work [49, 9, 18] also tried to build a better foundation model for medical images, by finetuning SAM on a large medical dataset and publishing the updated model weights. Some studies have also investigated SAM2 for medical images [12, 83, 50]. They found mixed results for 2D segmentation, improving over SAM for some modalities but with worse performance for others, and promising results for video segmentation [44].
Figure 1: a) Contribution overview: We finetune SAM on a large medical dataset to build MedicoSAM. We evaluate it for interactive and semantic segmentation. The latter requires training on additional annotated data (that could be generated via interactive segmentation), for 2D and 3D data. b) Results for interactive 2D segmentation, comparing MedicoSAM and other models derived from SAM. We report the average over 16 datasets for segmentation with a point (green) or box (yellow) prompt and segmentation after iterative correction starting from a point (dark green) or box (dark purple). c) Results for interactive 3D segmentation. We report the average over 6 different datasets for segmentation based on a single point or box. d, e). Results for semantic 2D and 3D segmentation. We report the average over 6 datasets in both cases. The three best methods are highlighted in decreasing shades of blue, darker indicates better results, gray otherwise.
@@
However, a comprehensive study that compares different approaches for improving SAM as a foundation model for medical images is so far missing. Specifically, no prior work has explored the impact of different finetuning strategies on the different criteria that a foundation model for medical segmentation should fulfill:
@@
    It should improve interactive segmentation. The imaging modalities and segmentation tasks in medicine are very diverse. Hence, a single model that can solve any medical segmentation task automatically is currently not feasible [^1] There exist some efforts to establish such models for specific modalities, most notably TotalSegmentator [69, 11] for CT and MRI images.. Better interactive segmentation will enable semi-automatic data annotation, leading to faster annotation times, either for data analysis or model training.
@@
    It should improve the performance for downstream tasks, in particular as a pretrained model for semantic segmentation. This would enable automating segmentation by supervised finetuning, potentially using annotations generated interactively.
@@
    It should be compatible with the original SAM library tools for data annotation (e.g. [46, 1]), enabling users to benefit from improved interactive segmentation.
Previous work has only addressed one of these aspects at a time: MedSAM [49] and SAMed2D [9] evaluate interactive segmentation (1), Gu et al. [18] study semantic segmentation (2). None of them explicitly study compatibility with user-friendly tools (3).
Our work closes this gap by comparing different training strategies on the dataset published by SAMed2d [9] and evaluating their effect on (1-3). Specifically, we evaluate own models trained with different strategies and published SAM-derived models on challenging medical segmentation tasks from four different categories: interactive 2D/3D segmentation and semantic 2D/3D segmentation. Where applicable, we also compare to SAM2 and MedSAM2 [50]. An overview of our approach is shown in Fig. 1a and a summary of our results in Fig. 1b-e. We find that domain specific finetuning clearly improves interactive segmentation in 2D and 3D, given the right training objective. For semantic segmentation, pretraining on medical data leads to improved results compared to the original SAM model, with competitive or better performance compared to nnU-Net [28] in 2D but worse performance in 3D. Our software and our best model, which we call MedicoSAM, are available at https://github.com/computational-cell-analytics/medico-sam.
@@ 2 Methods
We provide a summary of the contributions made by SAM [33], focusing on its training objective (Sec. 2.1). We then describe the finetuning strategies explored in our study (Sec. 2.2), including our contribution for pre-training semantic segmentation, our extension of SAM to interactive 3D segmentation (Sec. 2.3), our methods for 2D and 3D semantic segmentation (Sec. 2.4), and our evaluation methodology (Sec. 2.5).
2.1 Segment Anything Model
@@
SAM [33] is a vision foundation model for segmentation tasks. It consists of the image encoder, a ViT [13], the prompt encoder and the mask decoder. This architecture enables the model to solve interactive segmentation tasks based on user input, so called prompts. The image encoder processes the image and outputs an image representation. It contains the majority of parameters. SAM provides three different versions with different encoder sizes, ViT-Huge (ViT-h), ViT-Large (ViT-l) and ViT-Base (ViT-b). The prompt encoder processes the prompts, which can be point coordinates, either a positive point prompt (within the object of interest) or a negative point prompt (outside of the object of interest), a box coordinate, or a low-resolution mask. It outputs a representation of the prompts; point, box and mask prompts can be combined. The mask decoder processes the outputs of image encoder and prompt encoder to predict a mask of the object of interest and a score that estimates the prediction quality. It has two heads. One predicts a single mask and score, the other predicts three masks and scores. The second head is for the case of a single point prompt, which can result in ambiguities for part-object segmentation. Fig. 2a shows an overview of SAM’s architecture with additions made by us marked in orange.
@@
SAM was evaluated for a wide range of segmentation tasks in diverse image modalities, showing remarkable generalization. These capabilities are mainly due to two factors: its large and diverse training set and sophisticated training objective. The training dataset, called SA-1B, consists of 11 million images with 1 billion annotated objects. It was generated by human annotators using SAM for semi-automatic annotation, followed by retraining and further annotation with the updated model, repeated multiple times. The model was trained on this dataset using a supervised training objective that mimics interactive object annotation and correction: For a given ground-truth mask, the objective first samples either a point or box prompt and then corrects the model predictions with point prompts in multiple steps. In each step it computes $L_{mask}$, the loss between true and predicted mask as well as $L_{iou}$, the loss between the intersection over union (IOU) of true and predicted mask and the predicted score. These losses are accumulated and averaged at the end of a training iteration. See Alg. 1 for the pseudo-code of a training iteration. Besides this objective, the training proceeds as usual for deep neural networks by updating model weights with a version of stochastic gradient descent over multiple epochs.
Input: Images and target masks, hyperparameters $n_{obj}$, $n_{steps}$, $p_{box}$, $p_{mask}$, $e_{sem}$
@@
1 Sample minibatch of images and target masks
@@ 9     Compute bounding box of $m$, add as prompt to $p$
12     Sample random point from $m$, add as positive point prompt to $p$
@@ 23  Compute IOU $i$ between $\hat{m}$ and $m$
24  Compute regression loss $L_{iou}(\hat{i}$, ii)
@@ 27     Sample positive point from $m\,\&\,!\hat{m}$, add to $p$
28     Sample negative point from $m\,\&\,\hat{m}$, add to $p$
@@ 34    Run lines 12-21 with current $p$
37if $e_{sem}$ then
38  Compute binary target $b$ as union of all target masks
39  Predict binary mask $\hat{b}$ with additional segmentation decoder
40  Compute $L_{mask}(\hat{b},b)$ and add it to $L$
41
42Average losses in $L$, perform backprop
@@ 43 Update model parameters via optimizer
Algorithm 1 The training objective of SAM [33], according to [1]. We have added the optional joint pretraining of the segmentation decoder ($e_{sem}=1$).
Recently, SAM2 [61] has adapted this approach to interactive video segmentation by adding a memory bank that stores prompts and mask predictions from previous frames. SAM2 was trained on a large annotated video dataset. While SAM2 is promising in the medical domain to analyze videos or volumetric data, it has so far received fewer attention. Some studies [50, 12, 83] have evaluated it and, to our knowledge, only MedSAM2 [50] has provided an improved version of the model. Here, we include SAM2 and MedSAM2 in the evaluation for interactive segmentation, but do not finetune the model on medical data due to its recency and missing support in tools for medical data annotation.
@@ 2.2 Finetuning Segment Anything
We compare, implement, and extend different approaches for improving SAM for medical images by finetuning on a labeled dataset of medical images using a variation of Alg. 1. This training algorithm has not been published by SAM [33], so each publication has used a custom implementation. The finetuning strategies also differ in which model parts they update, and whether an adapter-based strategy like LoRA [25] is used.
@@
The authors of MedSAM [49] assemble a dataset of 1.5 million masks from CT, Endoscopy, MRI, X-Ray and other modalities, based on published data. They introduce a simple training objective that uses only box prompts, which are derived during training from the mask annotations. This objective corresponds to $n_{steps}=0,\,p_{box}=1,\,p_{mask}=0$ in Alg. 1. They use the ViT-b encoder and update all parameters of the image encoder and mask decoder, freezing the prompt encoder. The finetuned model is evaluated for interactive segmentation based on box prompts. The model is publicly available.
@@
The authors of SAM-Med2D [9] build the SA-Med2D-20M dataset [73], which consists of 20 million masks in 5 million images. The data covers ten different modalities (CT, Endoscopy, MRI, ultrasound, X-Ray and others) and is collected from published data. They finetune using with a training objective of $n_{steps}=8,\,p_{box}=0.5,\,p_{mask}=1$. They deviate from Alg. 1 in two ways: they sample either 1, 3, 5 or 9 points per step instead of a single positive and negative one (cf. lines 23-24) and they do not compute gradients for the prompt encoder in the steps corresponding to lines 22-29. They use ViT-b as image encoder and insert a low rank projection layer between attention and feed-forward layer of each transformer block, similar to [70]. Only these parameters of the image encoder are updated, others frozen. For prompt encoder and mask decoder all parameters are updated. In addition, they train the model with a smaller input image size of 256 x 256 pixels instead of 1024 x 1024 pixels used by SAM. They also study a simpler finetuning strategy called FT-SAM where only the mask decoder is updated during training, image and prompt encoder are frozen. The finetuned models are evaluated for interactive segmentation and are publicly available.
Figure 2: a) The SAM architecture for interactive segmentation consists of image encoder, prompt encoder (split into a part for mask prompts and for point/box prompts), and mask decoder. For 3D interactive segmentation, we propagate prompts across the depth axes. In addition, we add a convolutional decoder for automated segmentation (orange). This decoder is pre-trained with a binary segmentation task (blue masks), jointly with training for interactive segmentation.
b) Adaptation for 2D/3D semantic segmentation. A convolutional decoder predicts the segmentation output (same as ”Segmentation Decoder” in a)). For 3D segmentation, additional adapters are added to the image encoder and outputs of the segmentation decoder are computed per slice, stacked, and processed by a 3D convolution. Architecture modifications are highlighted in orange.
@@
The authors of [18] assemble a dataset comprising 100,000 masks in 300,000 images from CT, MRI, X-Ray and ultrasound. This dataset also contains unlabeled images used for self-supervised training. They compare two different finetuning objectives: updating the image encoder in a self-supervised manner using MAE [21] and using a simple supervised strategy with a box or a single point prompt, corresponding to $n_{steps}=0,\,p_{box}=0.5,\,p_{mask}=0$. They further compare the different model sizes (ViT-b, ViT-l, Vit-h) with and without the use of LoRA [25]. They evaluate these models for semantic segmentation tasks. None of the models are publicly available.
@@
  • MedSAM (adapted from [49]) uses only a box prompt, corresponding to $n_{steps}=0,\,p_{box}=1,\,p_{mask}=0$ in Alg. 1.
  • SimpleFT (adapted from [18]) uses a single box or a single point prompt, corresponding to $n_{steps}=0,\,p_{box}=0.5,\,p_{mask}=0$ in Alg. 1.
  • MedicoSAM uses the full objective with $n_{steps}=8,\,p_{box}=0.5,\,p_{mask}=0.5$ [^2] We have trained two different versions of this model, one with $p_{mask}=0.5$ and one with $p_{mask}=0$, see for details..
We set $n_{obj}=5$, use the Dice loss for $L_{mask}$ and the L2 loss for $L_{iou}$ in all cases. We also add an option to jointly pre-train the decoder for semantic segmentation. In this case, the decoder predicts a binary mask and we compute a loss between this prediction and the union of all target masks, corresponding to $e_{sem}=1$ in Alg. 1. See Sec. 2.4 for details on the decoder architecture. We finetune the models on SA-Med2D-20M [73]. We also benchmark the published models MedSAM [49], SAM-Med2D [9], FT-SAM [9], SAM [33], and where applicable SAM2 [61] and MedSAM2 [50]. We use ViT-b for all models, expect for MedSAM2, where only ViT-Tiny (ViT-t) is available.